# Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable emergent capabilities, transforming the execution of functional tasks by leveraging external tools for complex problems that require specialized processing or real-time data. While existing research expands LLMs access to diverse tools (e.g., program interpreters, search engines, weather/map apps), the necessity of using these tools is often overlooked, leading to indiscriminate tool invocation. This naive approach raises two key issues:(1) increased delays due to unnecessary tool calls, and (2) potential errors resulting from faulty interactions with external tools. In this paper, we introduce meta-cognition as a proxy for LLMs self-assessment of their capabilities, representing the model’s awareness of its own limitations. Based on this, we propose MeCo, an adaptive decision-making strategy for external tool use. MeCo quantifies metacognitive scores by capturing high-level cognitive signals in the representation space, guiding when to invoke tools. Notably, MeCo is fine-tuning-free and incurs minimal cost. Our experiments show that MeCo accurately detects LLMs’ internal cognitive signals and significantly improves tool-use decision-making across multiple base models and benchmarks.
@@
Wenjun Li*, Dexun Li*, Kuicai Dong, Cong Zhang, Hao Zhang, Weiwen Liu, Yasheng Wang, Ruiming Tang, and Yong Liu Huawei Noah’s Ark Lab {li.wenjun, lidexun, zhang.hao3, liuweiwen8, wangyasheng, tangruiming, liu.yong6@huawei.com}
**footnotetext: Both authors contributed equally to this work.
@@ 1 Introduction
Equipping large language models (LLMs) with tool-use capabilities allows them to overcome their limitations by accessing external/real-time data (Komeili, 2021; Tang et al., 2023), domain-specific knowledge (He-Yueya et al., 2023; Schick et al., 2024), and advanced specialized functionalities (Yang et al., 2023; Gao et al., 2023; Lu et al., 2024), thereby enabling them to handle more complex tasks beyond their inherent capabilities. While prior research has focused on expanding tool diversity (Qin et al., 2023; Hao et al., 2024) and optimizing their use (Patil et al., 2023; Shen et al., 2024), the decision-making process for determining when tools are necessary remains underexplored.
Indiscriminate invocation of external tools leads to two major issues: (1) increased latency (Qu et al., 2024; Wang et al., 2024), as external tools, such as search engine, typically operates slower compared to relying on internal knowledge retrieval, and (2) robustness risks, where dependence on external APIs increases the likelihood of errors due to tools malfunction or unnecessary tool-use (Qin et al., 2023; Lu et al., 2024; Wu et al., 2024).
@@
To address these limitations, we propose an adaptive tool-use strategy that improves decision-making in LLMs. Our approach, MeCo (Meta-Cognition-oriented trigger), enables LLMs to self-assess their capabilities and determine whether external tools are needed to address a user’s query. MeCo incorporates three key principles:
  • Meta-Cognition: We propose the concept of meta-cognition, which involves evaluating an LLM’s understanding of its capabilities based on its representation space. This self-awareness is crucial for minimizing unnecessary tool use.
  • Effective Strategy Utilization: We design an efficient strategy that leverages the quantified meta-cognition feedback to dynamically adjust tool use decisions, significantly improving decision accuracy compared to existing methods.
  • Generability: We empirically validate MeCo’s ability to generalize across a wide range of diverse scenarios. Additionally, we treat adaptive Retrieval-Augmented Generation (RAG) as a special case of tool use and demonstrate the superior effectiveness of MeCo in adaptive RAG compared to baseline methods.
Building on the Representation Engineering (RepE) framework (Zou et al., 2023), we develop a computationally efficient plug-in module to assess LLM meta-cognition. Our analysis reveals that meta-cognition produces strong signals, which can enhance interpretability and guide tool-use decisions. As shown in Figure 1, MeCo employs a dual-threshold policy to distinguish between strong and weak meta-cognitive signals, refining decision-making when uncertainty arises.
In summary, our contributions are four-fold: 1) We introduce the concept of adaptive tool use, which enhances both the efficiency and robustness of existing tool use paradigms in LLMs. 2) We integrate adaptive tool use and adaptive RAG within a unified framework, with their activation driven by a shared strategy based on meta-cognition detection. 3) We build a new benchmark, MeCa, to evaluate the effectiveness of our method. 4) We empirically demonstrate that MeCo significantly enhances the model’s awareness in tool utilization and RAG processes.
@@ 2 Background
Recent studies have explored LLMs’ internal representations to improve interpretability and understand their implicit beliefs (Bricken et al., 2023; Levinstein and Herrmann, 2024). Research (Zou et al., 2023; Liu et al., 2023a) shows that features like happiness, honesty, and confidence correspond to distinct, linearly separable directions in the representation space. Figure 2 illustrates the pipeline for training the corresponding feature probe. To detect these signals, contrastive instruction pairs can be used to induce their emergence.
Building on these insights, we can capture and control high-level functions $f$ (e.g., honesty) in model responses. We follow Zou et al. (2023) to design an experimental prompt $T^{+}_{f}$ that requires function execution and a reference prompt $T^{-}_{f}$ that does not. The instruction template is:
@@
where Rep represents the representation obtaining operation, $[-1]$ denotes the last token representation of $a^{k}$, and $A^{\pm}_{f}$ are the resulted activations consist of individual vectors.
Our goal is to learn a linear model to predict the direction of the function $A^{\pm}_{f}$ based on internal representations. Specifically, we apply PCA (Maćkiewicz and Ratajczak, 1993) in an unsupervised manner to pairwise difference vectors, deriving the first principal component $v_{f}$ (referred to as the probe) to identify function direction in the model’s responses. Equation 1 is applied at each layer of $M$ to derive layer-wise probes which are then used to interact with the LLM’s representations to monitor and control its behavior.
@@ 3 Approach
In the context of tool use, this involves assessing the model’s capabilities and limitations to determine whether a query can be answered independently or requires external tools, based on the query’s complexity and the sufficiency of the model’s internal knowledge.
To quantify meta-cognition, we train a probe that detects the model’s level of meta-cognitive awareness. This probe evaluates the rationale behind the model’s decision-making process, providing a score that reflects the model’s self-assessment accuracy. For instance, when the model receives a complex mathematical query, the meta-cognition probe assesses whether it correctly decides to solve the problem itself or delegate it to a calculator. A high meta-cognition score indicates accurate self-assessment, while a low score suggests either an incorrect attempt to solve it independently or unnecessary reliance on the tool.
@@ 3.1 Meta-Cognition Probe Extraction
Training a meta-cognition probe significantly differs from training probes for concepts like honesty or confidence. The latter concepts typically involve true-false statements about facts, such as "fire needs oxygen to burn" and "oxygen is harmful to human breathing". These statements are independent of user queries, meaning the model produces the same statements regardless of the user query.
In contrast, detecting the model’s internal cognition regarding tool use requires query-dependent responses. To achieve this, we employ leading proprietary LLM to generate user queries related to tool use and their corresponding responses (i.e., Yes/No responses with brief explanations). We then construct the training dataset following the procedures outlined in Section 2.
Notably, only a small dataset (of query-response pairs) suffices for strong probe performance. The analysis of the relationship between probe performance and the size of the training data is provided in Appendix D. Specifically, after collecting the instruction response pairs $(q_{i},a_{i})$, where $i$ denotes the index of the queries. We gather the sets of internal representations from the paired data and compute $A^{\pm}_{f}$ according to Eq. 1, and then apply PCA to the input $\{(-1)^{i}(A_{f,i}^{+}-A_{f,i}^{-})\}$ to obtain the first principal component $\nu_{f}$ as the meta-cognition probe. This principal component vector identifies the direction that represents the underlying meta-cognition concept.
After the training procedure described above, we obtain the meta-cognition probes across all model layers (e.g., 32 probes for Llama-3-8B). We compare our meta-cognition probe to existing proposed honesty Zou et al. (2023) and confidence probes Liu et al. (2024a) by evaluating the intermediate classification accuracy on held-out examples where the model is instructed to exhibit either honest/confident/strong meta-cognition or dishonest/lacking confidence/weak meta-cognition. The results of the probes are shown in Figure 3. As illustrated in Figure 3, our meta-cognition probe achieves near-optimal accuracy, significantly outperforming prior work.
@@
3.2 Decision-Making Strategy based on Meta-Cognition
With an accurate meta-cognition probe, we design a decision-making strategy based on detection results. For a given query, the LLM generates a response of $m$ tokens, each assigned a meta-cognition score across $n$ layers, forming a detection array of size $(m,n)$. The final decision—"Yes" (use external tools/RAG) or "No" (respond independently)—is derived from this array.
Reducing $m$ to 1.
We examine various prompting strategies (detailed in Appendix C) and find that the Yes/No+Explanation strategy, where the model answers with "Yes" or "No" followed by a brief explanation, yields the best performance. Therefore, we focus on the first token of the model’s response as it provides a clear signal of whether the model decides to rely on external tools. Extracting the meta-cognition score of the first token to represent the whole response simplifies our decision-making process, as calculating an overall meta-cognition score for the entire response is challenging due to varying response lengths and content across different queries. Since the model always responds with "Yes" or "No" as the first token, basing the trigger mechanism on the first token’s meta-cognition score is both reasonable and practical.
Reducing $n$ to 1.
In Zou et al. (2023) and Liu et al. (2024a), a mean score from multiple probes’ results is usually used to represent the token’s final quantification. However, our experiments show that scores predicted by different probes vary significantly, and simply averaging multiple scores does not yield accurate results. We found that probes in shallower layers (e.g., layer -5 to -2) tend to be more effective, with appropriate score distributions, ranges, and lower variances. Therefore, we select a single probe with the highest classification accuracy in the layer -5 to -2 (as shown in Figure 3) as the final predictor.
With meta-cognition scores distilled into a single scalar value, we apply the dual-thresholding strategy shown in Figure 1 to determine the optimal thresholds, $l_{yes}$ and $l_{no}$, using validation data. These thresholds are then applied to the test data.
@@ 4 Benchmark-MeCa
We evaluate MeCo using Metatool (Huang et al., 2023) and introduce a new benchmark, Meta-Cognitive Tool Assessment (MeCa), where each query undergoes human review. MeCa extends Metatool by incorporating a broader range of scenarios to assess tool usage and adaptive RAG.
Metatool comprises 1,040 queries designed to evaluate whether LLMs recognize when to use external tools. It includes 166 tools sourced from OpenAI’s plugin list (OpenAI, 2023). In Metatool, LLMs must decide on tool usage based solely on user queries, without tool names or descriptions. However, Metatool has limitations: 1) queries lack supplementary information or explicit tool provisions, whereas real-world tasks involve more complex intents and diverse requirements; 2) it primarily focuses on single-turn tool usage decisions. To address these gaps and ensure a more robust evaluation of MeCo, we developed MeCa, including two main components, MeCa-Tool and MeCa-RAG.
MeCa-Tool enhances Metatool by expanding tool-related assessments into three key categories:
  • Tool Usage Assessment: Evaluates whether an LLM should invoke external tools.
@@
  • Provided Tool Evaluation: Tests the LLM’s ability to determine tool usage when provided with a predefined set of tools.
@@
  • Multi-turn Interaction: Assesses tool usage decisions in multi-turn dialogues, requiring adaptation to evolving contexts.
@@
Specifically, we create six evaluation tasks in MeCa-Tool to systematically assess an LLM’s ability to make tool-related decisions across different scenarios. Tasks 1 and 4 evaluate whether an external tool is necessary to solve a query. Tasks 2 and 5 test the LLM’s ability to determine the relevance of a provided tool for a given query, including cases where the tool is irrelevant. Tasks 3 and 6 further extend this evaluation by presenting multiple tools (ranging from 2 to 5) and requiring the LLM to select the appropriate one. Notably, Tasks 1–3 focus on single-turn scenarios, while Tasks 4–6 incorporate multi-turn dialogues, assessing how well the LLM adapts its tool-use decisions in evolving conversational contexts. MeCa-Tool significantly expands Metatool by covering six tasks with 7,000 queries, providing a more diverse and comprehensive evaluation framework. The query composition of Metatool and MeCa-Tool is illustrated in Figure 4. Table 5 in Appendix A provides detailed statistics for each task.

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable emergent capabilities, transforming the execution of functional tasks by leveraging external tools for complex problems that require specialized processing or up-to-date data. While existing research expands LLMs access to diverse tools (e.g., program interpreters, search engines, calculators), the necessity of using these tools is often overlooked, leading to indiscriminate tool invocation. This naive approach raises two key issues: increased latency due to unnecessary tool calls, and potential errors resulting from faulty interactions with external tools. In this paper, we introduce meta-cognition as a proxy for LLMs self-assessment of their capabilities, reflecting the model’s awareness of its own limitations. Based on this, we propose MeCo, an adaptive decision-making strategy for external tool use. MeCo quantifies metacognitive scores by capturing high-level cognitive signals in the representation space, guiding when to invoke tools. Notably, MeCo is fine-tuning-free and incurs minimal cost. Experiments across multiple backbone models and benchmarks show that MeCo reliably detects LLMs’ internal cognitive signals and significantly improves tool-use decision-making.
@@
[…]
@@ 1 Introduction
Equipping large language models (LLMs) with tool use capabilities allows them to overcome their limitations by accessing external or up-to-date information (Komeili, 2021; Tang et al., 2023), domain-specific knowledge (He-Yueya et al., 2023; Schick et al., 2024), and advanced specialized functionalities (Yang et al., 2023; Gao et al., 2023; Lu et al., 2024), thereby enabling them to handle more complex tasks beyond their inherent abilities. While prior research has focused on expanding the tool arrays (Qin et al., 2023; Hao et al., 2024) and optimizing their use (Patil et al., 2023; Shen et al., 2024), the decision-making process for determining when tools are necessary remains underexplored.
Naively invoking tools indiscriminately leads to two major issues: (1) increased latency (Qu et al., 2024; Wang et al., 2024), as external tools, such as search engine, typically operates slower compared to relying on internal knowledge retrieval, and (2) robustness risks, where dependence on external APIs increases the likelihood of errors due to tools malfunction or unnecessary tool use (Qin et al., 2023; Lu et al., 2024; Wu et al., 2024).
@@
To address these limitations, we propose an adaptive tool-use strategy that improves decision-making in LLMs. Our approach, MeCo (Meta-Cognition-oriented trigger), enables LLMs to self-assess their capabilities and decide whether external tools are needed to answer a given query. MeCo incorporates three key principles:
  • Meta-Cognition: We define meta-cognition in the context of LLM tool useas the model’s ability to evaluate its own competence based on internal representations. This self-awareness is crucial for minimizing unnecessary tool use.
  • Effective Strategy Utilization: We design an efficient strategy that leverages the quantified meta-cognition signals to dynamically adjust tool use decisions, significantly improving decision accuracy compared to baseline methods.
  • Generability: We demonstrate that MeCo generalizes well across diverse tasks and domains. Additionally, we treat adaptive Retrieval-Augmented Generation (RAG) as a special case of tool use and show MeCo outperforms standard baselines.
Building on the Representation Engineering (RepE) (Zou et al., 2023), we develop a computationally efficient plug-in module to assess meta-cognition in LLMs. Our analysis reveals that meta-cognition provides strong and interpretable signals that can effectively guide tool use decisions. As illustrated in Figure 1, MeCo employs a dual-thresholding policy to distinguish between strong and weak meta-cognitive signals, refining decision-making when uncertainty arises.
In summary, our contributions are four-fold: 1) We introduce the concept of adaptive tool use, which enhances both the efficiency and robustness of LLM tool use paradigms. 2) We unify adaptive tool use and adaptive RAG under a shared framework, with their activation driven by meta-cognition-based decision-making. 3) We build a new benchmark, MeCa, to systematially evaluate the effectiveness of our approach. 4) We empirically demonstrate that MeCo significantly enhances model awareness in both tool utilization and RAG scenarios.
@@ 2 Background
Recent studies have explored the internal representations of LLMs to improve interpretability and understand their implicit beliefs (Bricken et al., 2023; Levinstein and Herrmann, 2024). Prior work (Zou et al., 2023; Liu et al., 2023a) shows that abstract features like happiness, honesty, and confidence correspond to distinct, linearly separable directions in the representation space. Figure 2 illustrates the pipeline for training these feature probes. To detect these signals, contrastive instruction pairs are typically used to induce their emergence.
Building on these findings, we can capture and manipulate high-level functions $f$ (e.g., honesty) in model responses. We follow Zou et al. (2023) to design an experimental prompt $T^{+}_{f}$ that requires function execution and a reference prompt $T^{-}_{f}$ that does not. The instruction template is as follows:
@@
where Rep denotes the representation obtaining operation, [1][-1] is the last token representation of $a^{k}$, and $A^{\pm}_{f}$ are the resulted activations consist of individual vectors.
The goal is to learn a linear model to predict the direction of the function $A^{\pm}_{f}$ based on internal representations. Specifically, we apply PCA (Maćkiewicz and Ratajczak, 1993) in an unsupervised manner to pairwise difference vectors, deriving the first principal component $v_{f}$ (referred to as the probe) to identify function directions in the representation space. Equation (1) is applied at each layer of $M$ to derive layer-wise probes which are then used to interact with the LLM’s representations to monitor and control its behavior.
@@ 3 Approach
In the context of tool use, this involves assessing the model’s capabilities and limitations to determine whether a query can be answered independently or requires external tools, based on the complexity of the query and the sufficiency of model’s internal knowledge.
To quantify meta-cognition, we train a probe that detects the model’s level of meta-cognitive awareness. This probe evaluates the rationale behind the model’s decision-making process, providing a score that reflects the model’s self-assessment accuracy. For instance, when the model receives a complex mathematical query, the meta-cognition probe assesses whether it correctly decides to solve the problem itself or delegate it to a calculator. A high meta-cognition score indicates accurate self-assessment, while a low score suggests uncertainty in the model’s decision.
@@ 3.1 Meta-Cognition Probe Extraction
Training a meta-cognition probe differs fundamentally from training honesty or confidence probes. The latter probes are typically trained on true/false factual statements, e.g., "fire needs oxygen to burn" and "oxygen is harmful to human breathing". These statements are independent of specific user queries, meaning the model produces consistent internal representations regardless of the prompt context.
In contrast, detecting internal cognitive signals related to tool use requires query-dependent responses. To achieve this, we use leading proprietary LLMs to generate user queries related to tool use and their corresponding responses (i.e., binary Yes/No answers with brief explanations). We then construct the training dataset following the procedures outlined in Section 2.
Notably, only a small dataset suffices for strong probe performance. Appendix D provides an analysis of the relationship between probe performance and training data size. Specifically, after collecting the instruction-response pairs $(q_{i},a_{i})$, we extract the model’s internal representations and compute the contrastive representation differences $A^{\pm}_{f}$ according to Equation (1). We then apply PCA to the set of difference vectors $\{(-1)^{i}(A_{f,i}^{+}-A_{f,i}^{-})\}$ to obtain the first principal component $\nu_{f}$ as the meta-cognition probe, which identifies the direction of the underlying meta-cognition signal.
After the above procedures, we obtain meta-cognition probes across all model layers (e.g., 32 probes for Llama-3-8B). We evaluate our meta-cognition probe against existing probes for honesty Zou et al. (2023) and confidence Liu et al. (2024a) by measuring the intermediate classification accuracy on held-out examples where the model is instructed to exhibit either honest/confident/strong meta-cognition or dishonest/lacking confidence/weak meta-cognition. Intermediate classification accuracy refers to the probe’s ability to correctly predict the directional alignment, i.e., whether the token representation points in a positive or negative direction along the concept axis (honesty/confidence/meta-cognition). This metric does not reflect final tool-use decision accuracy. The results shown in Figure 3 indicates our meta-cognition probe achieves near-optimal accuracy, significantly outperforming prior probes.
@@
3.2 Decision-Making Strategy based on Meta-Cognition
With an accurate meta-cognition probe, we design a decision-making strategy based on the detection results. For a given query, the LLM generates a response of $m$ tokens, each with a meta-cognition score across $n$ layers, forming a detection array of size $(m,n)$. The final decision—"Yes" (use external tools/RAG) or "No" (respond independently)—is derived from this array.
Reducing $m$ to 1.
We examine various prompting strategies (detailed in Appendix C) and find that the Yes/No+Explanation strategy, where the model answers with "Yes/No" followed by a brief explanation, yields the best performance. Hence, we focus on the first token of the model’s response, as it provides a clear signal of whether the model decides to rely on external tools. Extracting the meta-cognition score of the first token to represent the whole response simplifies our decision-making process, since calculating an overall meta-cognition score for the entire response is challenging due to varying response lengths and content across different queries. As the model consistently responds with "Yes/No" as the first token, basing the trigger mechanism on the first token’s meta-cognition score is both reasonable and practical.
Reducing $n$ to 1.
In Zou et al. (2023) and Liu et al. (2024a), a mean score from multiple probes’ results is typically used to represent the token’s final quantification. However, our experiments show that scores predicted by different probes vary significantly, and simply averaging them does not yield accurate results. We found that probes in shallower layers (e.g., layers -5 to -2) tend to be more effective, with appropriate score distributions, ranges, and lower variances. Therefore, we select the single probe with the highest classification accuracy within layers -5 to -2 (as shown in Figure 3) as the final predictor.
Dual-Thresholding.
After distilling the meta-cognition scores into a single scalar value, we adopt a simple yet effective dual-thresholding strategy (as illustrated in Figure 1) to determine the optimal thresholds, $l_{yes}$ and $l_{no}$, using validation data. These thresholds are then directly applied to the test data. As shown in Figure LABEL:fig:metacog_distribution, we empirically observe that correct Yes responses tend to have higher meta-cognition scores than incorrect Yes responses, while correct No responses tend to have lower scores than incorrect No responses. Notably, in both pre- and post-fine-tuning experiments, there is a clear gap between the meta-cognition scores of correct and incorrect responses. Our decision-making strategy identifies and leverages this gap to make better decisions. This pattern arises because the meta-cognition score is influenced by the token embedding of Yes/No, making scores between the two classes not directly comparable. Consequently, Yes scores should only be compared with other Yes scores, and similarly for No scores.
Driven by this empirical observation, we retain the model’s decision when the meta-cognition score $\nu$ is higher than $l_{yes}$ for Yes decisions, or lower than $l_{no}$ for No decisions; and override the model’s decision when $\nu$ is lower than $l_{yes}$ for Yes decisions, or higher than $l_{no}$ for No decisions. We demonstrate the robustness and effectiveness of this dual-thresholding strategy through comprehensive evaluations on both the MetaTool and MeCa benchmarks in Section 6.
@@ 4 Benchmark-MeCa
We evaluate MeCo using MetaTool (Huang et al., 2023) and introduce a new benchmark, Meta-Cognitive Tool Assessment (MeCa), in which each query undergoes human review. MeCa extends MetaTool by incorporating a broader range of scenarios to assess adaptive tool use and RAG.
MetaTool comprises 1,040 queries designed to evaluate whether LLMs recognize when to use external tools. In MetaTool, LLMs must decide on tool usage based solely on user queries, without tool names or descriptions. We identify the following limitations of MetaTool: 1) queries lack supplementary information or explicit tool provisions, whereas real-world tasks involve more complex intents and diverse requirements; 2) it primarily focuses on single-turn tool usage decisions.
To address these gaps and provide a more robust evaluation of MeCo, we introduce MeCa, which includes two main components: MeCa-Tool and MeCa-RAG. MeCa-Tool extends MetaTool by expanding tool-related assessments across three key categories:
  • Tool Use Assessment: Evaluates whether the LLM should invoke external tools.
@@
  • Provided Tool Evaluation: Tests the LLM’s ability to decide on tool usage when given a predefined set of tools.
@@
  • Multi-turn Interaction: Assesses tool use decisions in multi-turn dialogues that require adaptation to evolving contexts.
@@
Specifically, we create six evaluation tasks in MeCa-Tool to systematically assess an LLM’s ability to make tool-related decisions across various scenarios. Tasks 1 and 4 evaluate whether an external tool is necessary to solve a given query. Tasks 2 and 5 assess the LLM’s ability to determine the relevance of a provided tool, including cases where the tool is irrelevant. Tasks 3 and 6 further extend this evaluation by presenting multiple tools (ranging from 2 to 5) and requiring the LLM to select the appropriate one. Notably, Tasks 1–3 focus on single-turn settings, while Tasks 4–6 involve multi-turn dialogues, testing the LLM’s ability to adapt tool-use decisions to evolving conversational contexts. MeCa-Tool significantly expands MetaTool by covering six tasks with 7,000 queries, providing a more diverse and comprehensive evaluation framework. The query composition of MetaTool and MeCa-Tool is illustrated in Figure 9, and detailed task statistics are provided in Table 7 in Appendix A.
Beyond tool use, MeCa-RAG evaluates adaptive RAG—determining whether a query can be answered directly by the LLM or requires external retrieval. RAG is a special case of tool use, where the LLM’s internal knowledge is insufficient and necessitates using of a search engine to access external information. MeCa-RAG includes:
  • Positive RAG: Cases where retrieval is essential to answer complex queries or those requiring up-to-date information.
  • Negative RAG: Cases where the LLM can provide correct answers using its internal knowledge without retrieval.
To curate the MeCa-Tool dataset, we employed a meticulous approach that began with collecting diverse scenarios from various online corpora, ensuring our synthetic APIs and conversations were grounded in realistic contexts. Based on these scenarios, we designed 500 distinct synthetic APIs emulating real-world applications across various domains. We then generated queries by randomly sampling from this API pool—queries that either require tool invocation, rely on the LLM’s internal knowledge, or expose cases where the provided APIs are insufficient—followed by rigorous human verification for accuracy. For the MeCa-RAG dataset, we selected common fact-based data from the RepE (Zou et al., 2023) dataset (e.g., "The Earth orbits the Sun”) to generate negative queries that do not require retrieval, while positive queries were generated from crawled recent news unseen by LLMs, thereby necessitating retrieval of new information.
By incorporating both adaptive tool usage and adaptive RAG, MeCa serves as a robust benchmark for assessing LLM decision-making in complex scenarios. Detailed benchmark statistics and creation methodology for MeCa-RAG can be found in Appendix A.
5 Experiment Setup
Baselines: We evaluate MeCo against two baselines: Naive and $P_{\text{Yes}}$. The Naive baseline determines "Yes" or "No" based solely on the first token generated by the LLM, where "Yes" indicates the need for external tools, and "No" indicates otherwise. The $P_{\text{Yes}}$ baseline refines this by computing a Yes-score,
$\text{Yes-score}=\frac{P(\text{Yes }|\text{ Prompt})}{P(\text{Yes }|\text{ Prompt})+P(\text{No }|\text{ Prompt})}$
Backbone LLMs: We employ Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Llama-3-70B-Instruct as backbone models to evaluate MeCo. For brevity, we refer to them as LM3-8B, Mist-7B, and LM3-70B. Additionally, we fine-tune these models on a dataset of 4,000 tool-use query-response pairs generated by GPT-4-turbo, denoting the fine-tuned versions as LM3-8B-sft, Mist-7B-sft, and LM3-70B-sft. We include the 70B variant to demonstrate MeCo ’s effectiveness on large-scale models and its potential applicability in industry-grade deployments.
Evaluation: The primary evaluation metric is decision accuracy—whether the model correctly identifies when external tools/RAG are genuinely necessary. Additional metrics (precision, recall, etc.) are reported in Appendix C.
Prompting Strategies: We experimented with various prompting strategies, including "Yes/No" with and without explanations and Chain-of-Thought (CoT; Wei et al. (2022)). The best performance was achieved using the "Yes/No + Explanation" strategy, which is used throughout this paper. Full results for all prompting variants are provided in Appendix C. We further employ two types of prompts in our experiments: 1) prompts with context, which include specific reasons why LLMs require external tools to complete user tasks, plus five randomly sampled in-context examples to guide the model’s decisions; and 2) prompts without context, a concise version containing only the instruction and query. Exact prompt templates are detailed in Appendix D.
6 Experiments
We conduct extensive experiments to empirically reveal the effectiveness of MeCo on MetaTool and MeCa. Specifically, we assess MeCo in adaptive tool use on both MetaTool and MeCa-Tool and in adaptive RAG on MeCa-RAG.
6.1 Reduced Tool Invocations and Latency
Before delving into detailed benchmark performance, we first demonstrate that using MeCo in adaptive tool use and adaptive RAG reduces unnecessary tool invocations and overall latency. For this evaluation, we use the LM3-8B model.
We begin by evaluating the number of tool invocations on MetaTool, which contains an equal proportion of queries that require and do not require external tools. Table 1 summarizes the number of tool invocations made by each method. The Naive method invokes 296 tool calls before fine-tuning and 277 after fine-tuning. $P_{\text{Yes}}$ improves efficiency by reducing the number of invocations to 208 and 273, respectively. Notably, MeCo achieves the lowest number of tool invocations—206 before fine-tuning and 240 after—while maintaining the highest decision accuracy. These results highlight MeCo’s effectiveness in addressing the two key issues associated with indiscriminate tool use:
  • Reduced Latency: By minimizing unnecessary tool calls, MeCo lowers response latency, particularly for queries that do not require external tools (e.g., retrieval tools can take up to 14.8 seconds, as shown in Table 2).
  • Improved Robustness: Avoiding unnecessary tool invocations reduces reliance on potentially unreliable external APIs, thereby mitigating risks and malfunctions caused by unnecessary tool use.
Method Pre Fine-tuning Post Fine-tuning
Always Call 520 (50.0%) 520 (50.0%)
Naive 296 (61.9%) 277 (82.1%)
$P_{\text{Yes}}$ 208 (63.5%) 273 (81.7%)
MeCo 206 (65.0%) 240 (84.3%)
Table 1: Number of tool invocations and decision accuracy (in parentheses) of different methods on MetaTool.
Secondly, we evaluate the latency overhead introduced by MeCo on MeCa-RAG. This experiment adopts a standard RAG pipeline involving a retrieval tool, consisting of the following stages:
  • Pre-processing: Queries are rewritten to be clearer and more specific for search. If the query is complex, it is decomposed into simpler, related sub-queries.
  • Search Engine: Google Search is used to retrieve the top-5 most relevant web pages.
  • Post-processing: The model integrates the original query with the retrieved content to generate an answer. When needed, relevant details are summarized or extracted to directly address the user’s query.
Method Search Overhead (s) w/ ctx w/o ctx
Always Call 14.8 0.00 0.00
Naive 14.8 0.94 0.85
$P_{\text{Yes}}$ 14.8 0.98 0.89
MeCo 14.8 1.74 0.95
Table 2: Increased latency (in seconds) introduced by different methods when deciding whether to invoke a search on MeCa-RAG. With context and without context are abbreviated as w/ cts and w/o ctx for brevity.
The average response time for invoking the retrieval tool (including pre-processing, search, and post-processing) is 14.8 seconds. While MeCo introduces a slight latency overhead due to the computation of meta-cognition scores, this cost is minimal. As shown in Table 2, the additional latency incurred by MeCo is negligible relative to the overall retrieval time, particularly for searches without context. For instance, MeCo adds only 0.95 seconds (a 6.4% increase over $P_{\text{Yes}}$), which is marginal compared to the 14.8-second baseline. More importantly, MeCo avoids unnecessary tool invocations, particularly in multi-round retrieval settings, thereby improving overall response efficiency and mitigating the significant delays introduced by external tools.